# Code Interviews: Design and Evaluation of a More Authentic Assessment for Introductory Programming Assignments

### v1

@@ 5. Conclusion
We designed code interviews with the goal of creating a more authentic and robust way to evaluate standardized take-home programming assignments. The results of our action research paint a nuanced picture about the potential benefits and challenges around the use of code interviews for supporting assessment in introductory programming, revealing several tensions in the design of our code interview format. Code interviews (1) pushed students to discuss their work, motivating more nuanced but sometimes repetitive insights; (2) enabled peer learning, reducing stress in some ways but increasing stress in other ways; (3) scaled with TA-led sections, replacing familiar practice with an unfamiliar assessment; (4) focused on student contributions, limiting opportunities for TAs to give guidance and feedback.
As a primarily qualitative and exploratory experience report, our work raises further questions for investigation and opportunities to redesign code interviews to address these tensions. In its current form, code interviews during both winter and spring quarter had a high degree of structure and roleplaying that restricted the ways TAs and students could participate: TAs asked questions while students answered them. Future work could envision code interviews as collaborations between students and TAs, blending elements of learning and creativity with assessment. What might it look like for students to pose questions for themselves or their peers to explore? What real-world practices can we draw on to motivate creativity, collaboration, and self-expression in assessment?
@@
Our study also triple-staffed sections with TAs, which is not sustainable in the long-term at our institution without making significant adjustments to the distribution of TA responsibilities. Would it be possible for students to conduct peer code interviews whereby students interview each other? Could peer code interviews enable peer learning, reduce student stress, and scale weekly assessment without feeling repetitive? Could these peer code interviews still be accurate assessments of course learning outcomes? How should TAs moderate this process and prepare students to be effective interviewers? Future work could apply lessons from the literature on near-peer mentoring to inspire the design of peer code interviews.
During both winter quarter and spring quarter, code interviews supplemented traditional marking and evaluation of take-home programming assignments. But code interviews do not need to relegated to a sidecar. As generative artificial intelligence influences the landscape for programming skills, code interviews seem particularly capable of adapting to new demands and new kinds of authentic programming skills. During spring quarter, code interviews asked specific questions about each weekly assignment with emphasis on the particular skills and learning outcomes that the course staff deemed most important for each assignment. Speaking anecdotally, code interviews—and isomorphic code writing problems in particular—provided TAs a very close look at students’ problem solving process and programming fluencies. Code interviews tell us much more than whether the student got the question right or wrong. Future work can explore the potential for code interviews to serve as diagnostic or formative assessments that support student learning rather than (or in addition to) their potential as grading instruments.

### v2

@@ 5. Conclusion
As a primarily qualitative and exploratory experience report, our work raises questions for further study. Code interviews during both quarters had TAs asking questions and students answering them, but student and TA preferences about the specific structure and format were mixed. By spring quarter, we felt the most suitable structure and format depended on the particular assessments and particular learning objectives. But future work could envision code interviews as a more collaborative effort between students and TAs, blending elements of learning with assessment. For instance, allowing students to pose questions to peers or TAs could foster critical thinking and a deeper engagement with their code. Code interviews could be perceived as a more authentic task if students were given more space to engage their curiosity, creativity, collaboration, self-expression, and ways of knowing.
@@
Additionally, our approach required triple-staffing sections with TAs, which is not sustainable in the long term without significant adjustments to TA responsibilities. While this workload is likely lower than the alternative of offering (and then grading) a paper-and-pencil exam during the same time frame, the high frequency of code interviews poses a challenge. Should frequent assessment be desirable, consider peer code interviews whereby students interview each other. Could these peer code interviews enable peer learning, reduce student stress, and enable frequent assessment at scale without sacrificing accuracy? Future research could build on the literature around near-peer mentoring to guide the design of effective peer-led code interviews, exploring how TAs might best moderate this process and prepare students to act as interviewers.
During both winter and spring, code interviews supplemented traditional assessment of take-home programming assignments. However, code interviews could take a more central role in the assessment process, particularly as generative AI affects the importance of particular programming skills. In this context, code interviews have the potential to adapt to new demands by enabling authentic assessment of programming skills in a way that supports academic integrity. During spring quarter, code interviews focused on specific skills and learning outcomes that course staff prioritized, allowing TAs to observe students’ problem-solving processes and programming fluency more closely. Anecdotally, isomorphic code writing problems in particular provided TAs a very close look at students’ problem solving process and programming fluency.
While our results primarily focused on code interview formats, it is important to consider student demographics when designing assessments and course policies. About 40% of students in both quarters identified as English Language Learners, which could have influenced perceptions about the stressfulness of the format. In addition to arriving at an answer to the TA’s questions, students were also expected to clearly explain their answers too. Future work should explore the role of language as a mediating factor for code interviews. Additionally, the particular TA–student pairings could have implications for student comfort and performance. Are students who are interviewed by TAs who share the same social identity presentation as them more inclined to feel more comfortable? Although representation among course staff and among students may appear similar at a macro level, the specific pairings within large classes may exacerbate inequities.
Finally, future work could also explore the potential for code interviews to serve as more diagnostic or formative assessments by supporting student learning in real time and encouraging reflection on coding strategies. Moving forward, code interviews may valuable not only as grading instruments but also as tools for directly supporting student learning, offering detailed feedback to students about how they might improve their programming process.